Given the linear equation y = -6x – 2, find the y-coordinates of the points (-1, ), (-2, ), and (1, ). Please show all of your work. Plot these points and graph the linear equation. S
Answer by stanbon(75887)

Given the linear equation y = -6x – 2,
find the y-coordinates of the points
y = -6*-1-2
y = 6-2
y = 4
(-1,4),
-----------------
y = -6*-2-2
y = 12-2
y = 10
(-2,10),
-----------------
y = -6*1-2
y = -8
(1,-8).
====================================
Please show all of your work. Plot these points and graph the linear equation.
I'll leave that to you.
Cheers,
Stan H.